Solving the following equations, which shows the correct answer?
A) log10 (0.1) = 1
B) log16 2 = 4
C) log 3 9 = 2
D) log5125 = 25

I am stumped!

Answer by scott8148(5880)

a logarithm is an exponent that a base (b) is raised to, to get a value

if logb(a)=c then b^c=a

A) 10^1 = ?

B) 16^4 = ?

C) 3^2 = ?

D) 5^25 = ?